Order: Cefprozil 0.5 g.p.o q12h. Available :Cefprozil oral suspension usp 125mg per 5ml
Answer #1

Ans) Order: Cefprozil 0.5 g.p.o q12h.

Available :Cefprozil oral suspension usp 125mg per 5ml

0.5 × 1000 = 500 mg

125 mg = 5 ml

500 mg = x

x = 500 × 5 ÷ 125 = 20 ml